# Quantum Optimization for the Maximum Cut Problem on a Superconducting Quantum Computer

### v1

@@ Acknowledgements.
Table S4: Prefactor values for the runtime model of Eq. (S34) on the Rigetti QCSTM data of Fig. S18.
@@
We anticipate $c_{0}$ could be minimized by reducing network and service latency, or by in-service integration of the QRR* algorithm. We suspect $c_{N}$ might be eliminated by horizontally scaling classical compute resources so all modular parts of the control system are initialized in $O(1)$ time. $c_{N.p}$ could be improved by increasing the rate at which individual cards in the control system can be programmed, while $c_{N.n_{\textrm{ex}}}$ could be improved with faster measurement classification and higher bandwidth in returning bit strings to the QRR* algorithm. $c_{n_{\textrm{ex}}}$ resolves to the expected value accounting for a default passive reset delay ($200~{}\mu\text{s}$) and measurement operations, and could be improved by an active reset protocol [110]. Finally, $c_{N.p.n_{\textrm{ex}}}$ establishes the per-gate-layer cost, given the total number of layers in a linear-chain QAOA ansatz of depth $p$ is $Np$ [33]; the fitted value of $446$ nanoseconds is generally consistent with the timing of three ISWAP plus one-qubit gates in Tab. S1 (see also Sec. S4.1.2 on compilation), noting that this measurement was conducted across a different set of qubits and with the possibility of parallel one-qubit and two-qubit operations.
@@
Assuming adoption of an active reset protocol [110] and elimination of all other overheads tabulated above, we establish the quantum computer runtime in our analyses of Fig. 2 of the main text as the current circuit duration for the performance data of Tab. S1. Precisely
$\tilde{t}_{\textrm{ex}}=n_{\textrm{ex}}\Bigl{[}t_{\textrm{init}}+p\bigl{(}t_{\textrm{mx}}+Nt_{\textrm{ps}}\bigr{)}+t_{\textrm{mes+res}}\Bigr{]},$ (S35)
@@
where $t_{\textrm{init}}$ accounts for the initial Hadamard gate, $t_{\textrm{mx}}$ the mixer time, $t_{\textrm{ps}}$ the phase separator time based on a linear architecture topology requiring a swap network [33], and $t_{\textrm{mes+res}}$ the measurement time and active reset of the qubit before the next run. The compilation overhead of the QAOA circuit into hardware-native gates (see Sec.  S4.1.2) leads to
$t_{\textrm{init}}=t_{\textrm{1Q}},\quad t_{\textrm{mx}}=2t_{\textrm{1Q}},\quad t_{\textrm{ps}}=4t_{\textrm{1Q}}+3t_{\textrm{2Q}},$ (S36)
@@
where $t_{\textrm{1Q}}\equiv t_{\texttt{Rx}}$ and $t_{\textrm{2Q}}\equiv t_{\texttt{ISWAP}}$ are the one- and two-qubit gate durations, respectively. The measurement time includes reading the qubit ($\simeq 1\mu\text{s}$) and performing an active reset based on its measured value, which involves applying a one-qubit gate and a classical feedback-loop limited by network latency of $\simeq 1~{}\mu\text{s}$. This reading/active reset strategy is repeated three times in a row, resulting in $t_{\textrm{mes+res}}\simeq 6~{}\mu\text{s}$ [110]. Current average values for the one- and two-qubit gates operations are reported in Tab. S1 where $t_{\texttt{Rx}}=40$ ns and $t_{\texttt{ISWAP}}=122$ ns. This model, shown as the solid line in Fig. S18, is approximately three times faster than the asymptotic performance measured on Rigetti QCSTM.
@@
A further threefold reduction in quantum computer runtime could be achieved by improving gate speeds towards the state-of-the-art for transmon-based superconducting quantum processors [101]; this is shown as the dashed line in Fig. S18 where $t_{\texttt{Rx}}=20$ ns and $t_{\texttt{ISWAP}}=50$ ns are used. This model was used in Fig. 3 of the main text.
The Light Cone Technique on the Quantum Processor
@@
The light cone technique of Sec. S3.2 trades the simulation of a unique large QAOA circuit for many smaller ones. On a chip of $M$ qubits, one may be able to fit several such subcircuits, thus reducing the overall runtime burden of the light cone decomposition. From the data of Fig. S4, we can estimate the average size of the subcircuits for different values of $(N,p)$. From the data of Fig. S6, we can estimate the average number of such subcircuits for a given problem as a function of $(N,p)$. Without relying on graph isomorphism the average number of such subcircuits is $O(N)$; see Sec. S3.3. We do not rely on graph isomorphism for estimating runtimes, given its poor algorithmic scaling.
@@
To establish the runtime of the quantum relax-and-round algorithm, we consider the number of times the $M$-qubit quantum computer needs to be used to compute all nonzero elements of the correlation matrix, given an average number of subcircuits of average size. The depth of those average-size subcircuits is related to using a complete swap network embedding of the phase separator component of the QAOA circuit onto a linear hardware-native topology (see Sec. S4.1.1) where no optimal embedding is sought [33]. For example, if one has an $M=84$-qubit quantum computer and needs to execute $180$ different subcircuits involving on average $7$ qubits, $12$ such subcircuits can be executed in parallel across the full quantum computer. This operation will need to be repeated $15$ times to obtain results for all subcircuits.
Runtime on the Classical Processor
Figure S19: (a) Runtime for computing the correlation matrix made of $n_{\textrm{entries}}$ nonzero entries from $n_{\textrm{ex}}$ bit strings. The runtime is linear per entry and bit string, with the prefactor approximately $3.859(3)\times 10^{-10}$ per least-squares fitting. (b) Runtime for the eigendecomposition of the $N\times N$ adjacency ($\mathsf{W}$) and correlation matrices ($\mathsf{Z}$) at $p=1$ and $p=2$ using the Lanczos algorithm. Each data point is averaged over $20$ matrices. Units in both panels: seconds.
@@
At this stage, we assume we have $n_{\textrm{ex}}$ bit strings $\boldsymbol{b}=\{0,1\}^{N}$, each of length $N$. First, we loop over all non-trivially zero entries $n_{\textrm{entries}}$ of the upper triangle of the $N\times N$ correlation matrix $\mathsf{Z}$. For each such entry $\mathsf{Z}_{ij}=(\delta_{ij}-1)\langle\hat{Z}_{i}\hat{Z}_{j}\rangle$, we estimate the expectation value
@@
$\bigl{\langle}\hat{Z}_{i}\hat{Z}_{j}\bigr{\rangle}\approx\frac{4}{n_{\textrm{ex}}}\sum\nolimits_{k=1}^{n_{\textrm{ex}}}\left(b^{(k)}_{i}-\frac{1}{2}\right)\left(b^{(k)}_{j}-\frac{1}{2}\right),$ (S37)
@@
where $b^{(k)}_{i}$ is the value of bit $i$ in the $k$th bit string $\boldsymbol{b}^{(k)}$ from the $k$th quantum circuit execution. We report in Fig. S19a the runtime of a C++ implementation of the above strategy on a $64$Gb MacBook Pro with an Apple M1 Max chip. We find that the runtime per bit string and entry is linear and can be fitted to $3.859(3)\times 10^{-10}\times n_{\textrm{entries}}\times n_{\textrm{ex}}$.
@@
Once the correlation matrix $\mathsf{Z}$ has been constructed, the next step is to compute its leading eigenvectors and sign-round them. We employ a C++ implementation of the Lanczos method to find the $k=8$ leading eigenvectors of the correlation matrix. Each of the eigenvectors is then sign-rounded and their corresponding objective values computed. The one with the best objective value is returned as the solution of the quantum relax-and-round algorithm. The runtime on a $64$Gb MacBook Pro with an Apple M1 Max chip is reported in Fig. S19b for the adjacency ($\mathsf{W}$) and correlation matrices ($\mathsf{Z}$) at $p=1$ and $p=2$ with different sizes $N$. There is no clear dependence on the number of entries in the matrix ($\mathsf{W}$ is sparser than $\mathsf{Z}^{(p=1)}$, itself sparser than $\mathsf{Z}^{(p=2)}$). We attribute that to the faster convergence of Lanczos with better-conditioned matrices. When needed, we will assume that the eigendecomposition for the correlation matrix at $p=3$ has the same runtime as for $p=2$.
Tabulated Runtime Summary for Quantum Solver
@@
The average runtime of the greedy-enhanced quantum relax-and-round solver (QRR*) for different scenarios and number of variables $N$ are in Tab. S5. We report $t_{\textrm{Q}}$ and $t_{\textrm{C}}$, which are the runtimes of the quantum (based on Eq. (S35)) and the classical components, respectively. These values were used for generating Figs.2 and 3 in the main text.
@@
Table S5: Average runtime in milliseconds (ms) of the greedy-enhanced quantum relax-and-round solver (QRR*) for different numbers of variables $N$ on random $3$-regular graphs. The different columns correspond to the various scenarios explored. $t_{\textrm{Q}}$ refers to the quantum computing runtime (based on Eq. (S35)) on an $M$-qubit quantum computer. $t_{\textrm{C}}$ refers to the classical component of the runtime, consisting of the time to compute expectation values, perform the eigendecomposition and $10$ local greedy steps.
S5.9.2 Classical Solvers
Figure S20: Runtime in seconds on a $64$Gb MacBook Pro with an Apple M1 Max chip for collecting one sample using simulated annealing with $k$ sweeps on problems of size $N$. The dashed line shows an asymptotic linear relationship yields an average runtime of $1.79(2)\times 10^{-8}$ seconds per sweep and per variable.
We focus on classical solvers which are most competitive against the low-depth quantum relax-and-round algorithm considered in this work: Simulated annealing (Sec. S2.1.2) and the heuristic by Burer and Monteiro (Sec. S2.1.4). The time required for constructing the problem or loading into the solver is excluded from the runtime. From there, each of these algorithms has a runtime dictated by a control parameter provided as input: the number of sweeps for simulated annealing, and the runtime limit for the MQLib/Burer2002 implementation of the Burer-Monteiro heuristic.
For simulated annealing, we control the number of sweeps $k$, which is proportional to the total runtime. As described in Sec. S2.1.2, each sweep involves an update over $N$ variables, each having only $O(1)$ nearest neighbors. Thus, the runtime is also proportional to $N$. In Fig. S20, we show the runtime (measured on a $64$Gb MacBook Pro with an Apple M1 Max chip) for collecting one sample as a function of the problem size $N$ multiplied by the number of sweeps $k$. Using a linear asymptotic fit we estimate the average runtime per sweep and per variable to be $1.79(2)\times 10^{-8}$ seconds.
S5.10 Optimal Runtime to Match the Performance of the Greedy-Enhanced Quantum Relax-and-Round Solver
@@
$*t^{*}=t\bigr{/}P\Bigl{(}C(\boldsymbol{z}_{\textrm{cl}})\geq C(\boldsymbol{z}_{\textrm{QRR*}})\Bigr{)},$ (S38)
@@
where $t$ is the actual runtime of the classical solver and $P$ the probability that the classical solver returns a solution $\boldsymbol{z}_{\textrm{cl}}$ at least matching that of the quantum algorithm $*\boldsymbol{z}_{\textrm{QRR*}}$. For the solvers considered here, $P$ is intrinsically related to $t$ given that a longer runtime is more likely to return a better solution, and conversely, a smaller runtime $t$ is more likely to result in a lower value of $P$. Therefore, we seek the optimal time $t^{*}_{\textrm{opt}}$ for the classical solver to find a solution matching the quality of the QRR* solution, which concurrently minimizes $t$ and maximizes $P$. We consider three classical solvers: Gurobi, simulated annealing, and the Burer-Monteiro approach.
S5.10.1 Simulated Annealing and the Burer-Monteiro Solvers
@@
The runtime $t$ of simulated annealing in Eq. (S38) is proportional to the number of variables in the problem and the number of sweeps performed, as shown in Fig. S20. The Burer-Monteiro implementation in MQLib/Burer2002, which was used in this work, takes $t$ as an input parameter.
@@
With knowledge of the solution $*\boldsymbol{z}_{\textrm{QRR*}}$ for a given problem instance, we run both solvers between $10^{2}$ and $10^{4}$ times with random initial seeds for a fixed set of parameters (number of sweeps for simulated annealing and maximum runtime for MQLib/Burer2002) and collect a solution from each of these runs to estimate $P$. This operation is repeated over $1,000$ randomly generated problem instances for each size $N$. We then compute an average time $\mathbb{E}[t^{*}]$ over all the problem instances, and subsets of the hardest instances. The time $t^{*}$ is used to identify the hardest instances.
@@
The results for simulated annealing and MQLib/Burer2002 are shown in Figs. S21 and S22, respectively. The optimal time $t^{*}_{\textrm{opt}}$ depends on $N$, ranging from a fraction of milliseconds $(N=32)$ to a fraction of seconds ($N=4,096$). In the main text, we compare this optimal runtime with the runtime of the QRR* algorithm.
@@
Figure S21: Runtime required for simulated annealing to match the performance of the greedy-enhanced quantum relax-and-round solver as a function of the number of sweeps. We consider all problem instances ($100\%$), as well as the hardest $10\%$ and $1\%$. Each panel corresponds to different problem sizes: (a) $N=32$, (b) $N=64$, (c) $N=128$, (d) $N=256$, (e) $N=512$, (f) $N=1024$, (g) $N=2048$, (h) $N=4096$. Error bars indicate one standard deviation. Runtimes were measured on a $64$Gb MacBook Pro with an Apple M1 Max chip. Units: seconds.
@@
Figure S22: Runtime required for the MQLib/Burer2002 implementation of the Burer-Monteiro heuristic to match the performance of the greedy-enhanced quantum relax-and-round solver as a function of the runtime input parameter. We consider all problem instances ($100\%$), as well as the hardest $10\%$ and $1\%$. Each panel corresponds to different problem sizes: (a) $N=32$, (b) $N=64$, (c) $N=128$, (d) $N=256$, (e) $N=512$, (f) $N=1024$, (g) $N=2048$, (h) $N=4096$. Error bars indicate one standard deviation. Runtimes were measured on a $64$Gb MacBook Pro with an Apple M1 Max chip. Units: seconds.
S5.10.2 Gurobi Solver

### v2

@@ E.8.1 Run-time of the Greedy-Enhanced Quantum Relax-and-Round Solver
Table 4: Prefactor values for the run-time model of Eq. (36) (in seconds) on the Rigetti QCSTM data of Fig. 17.
@@
We anticipate $c_{0}$ could be minimized by reducing network and service latency, or by in-service integration of the QRR* algorithm. We suspect $c_{N}$ might be eliminated by horizontally scaling classical compute resources so all modular parts of the control system are initialized in $O(1)$ time. $c_{N.p}$ could be improved by increasing the rate at which individual cards in the control system can be programmed, while $c_{N.n_{\textrm{ex}}}$ could be improved with faster measurement classification and higher bandwidth in returning bit strings to the QRR* algorithm. $c_{n_{\textrm{ex}}}$ resolves to the expected value accounting for a default passive reset delay ($200~{}\mu\text{s}$) and measurement operations, and could be improved by an active reset protocol [108]. Finally, $c_{N.p.n_{\textrm{ex}}}$ establishes the per-gate-layer cost, given the total number of layers in a linear-chain QAOA ansatz of depth $p$ is $Np$ [35]; the fitted value of $446$ nanoseconds is generally consistent with the timing of three ISWAP plus one-qubit gates (see App. D.1.2), noting that this measurement was conducted across a different set of qubits and with the possibility of parallel one-qubit and two-qubit operations.
@@
Assuming adoption of an active reset protocol [108] and elimination of all other overheads tabulated above, we establish the quantum computer run-time in our analyses of Fig. 2 of the main text as the current circuit duration for the performance data of Tab. 1. Precisely
$\tilde{t}_{\textrm{ex}}=n_{\textrm{ex}}\Bigl{[}t_{\textrm{init}}+p\bigl{(}t_{\textrm{mx}}+Nt_{\textrm{ps}}\bigr{)}+t_{\textrm{mes+res}}\Bigr{]},$ (37)
@@
where $t_{\textrm{init}}$ accounts for the initial Hadamard gate, $t_{\textrm{mx}}$ the mixer time, $t_{\textrm{ps}}$ the phase separator time based on a linear architecture topology requiring a swap network [35], and $t_{\textrm{mes+res}}$ the measurement time and active reset of the qubit before the next run. The compilation overhead of the QAOA circuit into hardware-native gates (see App. D.1.2) leads to
$t_{\textrm{init}}=t_{\textrm{1Q}},\quad t_{\textrm{mx}}=2t_{\textrm{1Q}},\quad t_{\textrm{ps}}=4t_{\textrm{1Q}}+3t_{\textrm{2Q}},$ (38)
@@
where $t_{\textrm{1Q}}\equiv t_{\texttt{Rx}}$ and $t_{\textrm{2Q}}\equiv t_{\texttt{ISWAP}}$ are the one- and two-qubit gate durations, respectively. The measurement time includes reading the qubit ($\simeq 1\mu\text{s}$) and performing an active reset based on its measured value, which involves applying a one-qubit gate and a classical feedback-loop limited by network latency of $\simeq 1~{}\mu\text{s}$. This reading/active reset strategy is repeated three times in a row, resulting in $t_{\textrm{mes+res}}\simeq 6~{}\mu\text{s}$ [108]. Current average values for the one- and two-qubit gates operations are reported in Tab. 1 where $t_{\texttt{Rx}}=40$ ns and $t_{\texttt{ISWAP}}=122$ ns. This model, shown as the solid line in Fig. 17, is approximately three times faster than the asymptotic performance measured on Rigetti QCSTM.
@@
A further threefold reduction in quantum computer run-time could be achieved by improving gate speeds toward the state-of-the-art for transmon-based superconducting quantum processors [99]; this is shown as the dashed line in Fig. 17 where $t_{\texttt{Rx}}=20$ ns and $t_{\texttt{ISWAP}}=50$ ns are used. This model was used in Fig. 3 of the main text.
The Light Cone Technique on the Quantum Processor—
@@
The light cone technique trades the simulation of a unique large QAOA circuit for many smaller ones. On a chip of $M$ qubits, one may be able to fit several such subcircuits, thus reducing the overall run-time burden of the light cone decomposition. From the data of App. C.2, we can estimate the average size of the subcircuits for different values of $(N,p)$ and the average number of such subcircuits for a given problem as a function of $(N,p)$. Without relying on graph isomorphism the average number of such subcircuits is $O(N)$. We do not rely on graph isomorphism for estimating run-times, given its poor algorithmic scaling.
@@
To establish the run-time of the quantum relax-and-round algorithm, we consider the number of times the $M$-qubit quantum computer needs to be used to compute all nonzero elements of the correlation matrix, given an average number of subcircuits of average size. The depth of those average-size subcircuits is related to using a complete swap network embedding of the phase separator component of the QAOA circuit onto a linear hardware-native topology (see App. D.1.1) where no optimal embedding is sought [35]. For example, if one has an $M=84$-qubit quantum computer and needs to execute $180$ different subcircuits involving on average $7$ qubits, $12$ such subcircuits can be executed in parallel across the full quantum computer. This operation will need to be repeated $15$ times to obtain results for all subcircuits.
Run-time on the Classical Processor—
@@
At this stage, we assume we have $n_{\textrm{ex}}$ bit strings $\boldsymbol{b}=\{0,1\}^{N}$, each of length $N$. First, we loop over all nontrivially zero entries $n_{\textrm{entries}}$ of the upper triangle of the $N\times N$ correlation matrix $\mathsf{Z}$. For each such entry $\mathsf{Z}_{ij}=(\delta_{ij}-1)\langle\hat{Z}_{i}\hat{Z}_{j}\rangle$, we estimate the expectation value
@@
$\bigl{\langle}\hat{Z}_{i}\hat{Z}_{j}\bigr{\rangle}\approx\frac{4}{n_{\textrm{ex}}}\sum\nolimits_{k=1}^{n_{\textrm{ex}}}\left(b^{(k)}_{i}-\frac{1}{2}\right)\left(b^{(k)}_{j}-\frac{1}{2}\right),$ (39)
@@
where $b^{(k)}_{i}$ is the value of bit $i$ in the $k$th bit string $\boldsymbol{b}^{(k)}$ from the $k$th quantum circuit execution. We report in Fig. 22a the run-time of a C++ implementation of the above strategy on a $64$Gb MacBook Pro with an Apple M1 Max chip. We find that the run-time per bit string and entry is linear and can be fitted to $3.859(3)\times 10^{-10}\times n_{\textrm{entries}}\times n_{\textrm{ex}}$.
@@
Once the correlation matrix $\mathsf{Z}$ has been constructed, the next step is to compute its leading eigenvectors and sign-round them. We employ a C++ implementation of the Lanczos method to find the $k=8$ leading eigenvectors of the correlation matrix. Each of the eigenvectors is then sign-rounded and their corresponding objective values computed. The one with the best objective value is returned as the solution of the quantum relax-and-round algorithm. The run-time on a $64$Gb MacBook Pro with an Apple M1 Max chip is reported in Fig. 22b for the adjacency ($\mathsf{W}$) and correlation matrices ($\mathsf{Z}$) at $p=1$ and $p=2$ with different sizes $N$. There is no clear dependence on the number of entries in the matrix ($\mathsf{W}$ is sparser than $\mathsf{Z}^{(p=1)}$, itself sparser than $\mathsf{Z}^{(p=2)}$). We attribute that to the faster convergence of Lanczos with better-conditioned matrices. When needed, we will assume that the eigendecomposition for the correlation matrix at $p=3$ has the same run-time as for $p=2$.
Tabulated Run-time Summary for Quantum Solver—
@@
The average run-time of the greedy-enhanced quantum relax-and-round solver (QRR*) for different scenarios and number of variables $N$ are in Tab. 5. We report $t_{\textrm{Q}}$ and $t_{\textrm{C}}$, which are the run-times of the quantum and the classical components, respectively. These values were used for generating Figs. 2 and 3 in the main text. The quantum run-time is based on
$t_{\textrm{Q}}=$ $\mathbb{E}\left[{\tilde{t}_{\textrm{ex}}(N,p)}\right]$
×$\times\left\lceil\mathbb{E}\left[{n_{\textrm{subgraphs}}(N,p)}\right]\frac{\mathbb{E}\left[{n_{\textrm{qubits}}(N,p)}\right]}{M}\right\rceil,$ (40)
where $\mathbb{E}[\cdot]$ denotes the ensemble average over the subgraphs (light-cone induced subcircuits), $\tilde{t}_{\textrm{ex}}$ the circuit duration based on Eq. (37), $n_{\textrm{subgraphs}}$ the number of subgraphs (App. C.3), $n_{\textrm{qubits}}$ the number of qubits for encoding a subgraph. These numbers depend on the number of variables $N$ of the problem and the number of QAOA layers $p$. $M$ denotes the number of qubits available on the quantum computer.
@@
Table 5: Average run-time in milliseconds (ms) of the greedy-enhanced quantum relax-and-round solver (QRR*) for different numbers of variables $N$ on random $3$-regular graphs. The different columns correspond to the various scenarios explored. $t_{\textrm{Q}}$ refers to the quantum computing run-time (based on Eq. (40)) on an $M$-qubit quantum computer. $t_{\textrm{C}}$ refers to the classical component of the run-time, consisting of the time to compute expectation values, perform the eigendecomposition and $10$ local greedy steps.
E.8.2 Measuring the Run-time of Classical Solvers
Figure 23: Run-time in seconds on a $64$Gb MacBook Pro with an Apple M1 Max chip for collecting one sample using simulated annealing with $k$ sweeps on problems of size $N$. The dashed line shows an asymptotic linear relationship yields an average run-time of $1.79(2)\times 10^{-8}$ seconds per sweep and per variable.
We focus on classical solvers which are most competitive against the low-depth quantum relax-and-round algorithm considered in this work: Simulated annealing (App. B.1.2) and the heuristic by Burer and Monteiro (App. B.1.4). The time required for constructing the problem or loading into the solver is excluded from the run-time. From there, each of these algorithms has a run-time dictated by a control parameter provided as input: the number of sweeps for simulated annealing, and the run-time limit for the MQLib/Burer2002 implementation of the Burer-Monteiro heuristic.
For simulated annealing, we control the number of sweeps $k$, which is proportional to the total run-time. As described in App. B.1.2, each sweep involves an update over $N$ variables, each having only $O(1)$ nearest neighbors. Thus, the run-time is also proportional to $N$. In Fig. 23, we show the run-time (measured on a $64$Gb MacBook Pro with an Apple M1 Max chip) for collecting one sample as a function of the problem size $N$ multiplied by the number of sweeps $k$. Using a linear asymptotic fit we estimate the average run-time per sweep and per variable to be $1.79(2)\times 10^{-8}$ seconds.
E.9 Optimal Run-time to Match the Performance of the Greedy-Enhanced Quantum Relax-and-Round Solver
@@
where $t$ is the actual run-time of the classical solver and $P$ the probability that the classical solver returns a solution $\boldsymbol{z}_{\textrm{cl}}$ at least matching that of the quantum algorithm $*\boldsymbol{z}_{\textrm{QRR*}}$. For the solvers considered here, $P$ is intrinsically related to $t$ given that a longer run-time is more likely to return a better solution, and conversely, a smaller run-time $t$ is more likely to result in a lower value of $P$. Therefore, we seek the optimal time $t^{*}_{\textrm{opt}}$ for the classical solver to find a solution matching the quality of the QRR* solution, which concurrently minimizes $t$ and maximizes $P$. We consider three classical solvers: Gurobi, simulated annealing, and the Burer-Monteiro approach.
E.9.1 Simulated Annealing and the Burer-Monteiro Solvers
@@
The run-time $t$ of simulated annealing in Eq. (41) is proportional to the number of variables in the problem and the number of sweeps performed, as shown in Fig. 23. The Burer-Monteiro implementation in MQLib/Burer2002, which was used in this work, takes $t$ as an input parameter.
@@
With knowledge of the solution $*\boldsymbol{z}_{\textrm{QRR*}}$ for a given problem instance, we run both solvers between $10^{2}$ and $10^{4}$ times with random initial seeds for a fixed set of parameters (number of sweeps for simulated annealing and maximum run-time for MQLib/Burer2002) and collect a solution from each of these runs to estimate $P$. This operation is repeated over $1,000$ randomly generated problem instances for each size $N$. We then compute an average time $\mathbb{E}[t^{*}]$ over all the problem instances, and subsets of the hardest instances. The time $t^{*}$ is used to identify the hardest instances.
@@
The results for simulated annealing and MQLib/Burer2002 are shown in Figs. 24 and 25, respectively. The optimal time $t^{*}_{\textrm{opt}}$ depends on $N$, ranging from a fraction of milliseconds $(N=32)$ to a fraction of seconds ($N=4,096$). In the main text, we compare this optimal run-time with the run-time of the QRR* algorithm.
@@
Figure 24: Run-time required for simulated annealing to match the performance of the greedy-enhanced quantum relax-and-round solver as a function of the number of sweeps. We consider all problem instances ($100\%$), as well as the hardest $10\%$ and $1\%$. Each panel corresponds to different problem sizes: (a) $N=32$, (b) $N=64$, (c) $N=128$, (d) $N=256$, (e) $N=512$, (f) $N=1024$, (g) $N=2048$, (h) $N=4096$. Error bars indicate one standard deviation. Run-times were measured on a $64$Gb MacBook Pro with an Apple M1 Max chip. Units: seconds.
@@
Figure 25: Run-time required for the MQLib/Burer2002 implementation of the Burer-Monteiro heuristic to match the performance of the greedy-enhanced quantum relax-and-round solver as a function of the run-time input parameter. We consider all problem instances ($100\%$), as well as the hardest $10\%$ and $1\%$. Each panel corresponds to different problem sizes: (a) $N=32$, (b) $N=64$, (c) $N=128$, (d) $N=256$, (e) $N=512$, (f) $N=1024$, (g) $N=2048$, (h) $N=4096$. Error bars indicate one standard deviation. Run-times were measured on a $64$Gb MacBook Pro with an Apple M1 Max chip. Units: seconds.
E.9.2 Gurobi Solver